Nickel Inc. bought $100,000 of 3-year, 6% bonds as an investment on December 31, 2017 for $106,000. The investment receives interest annually and Nickel uses straight-line amortization. On May 1, 2018, the issuer retired $10,000 of the bonds at 110. As a result of the retirement, Nickel will report a:

a. $467 gain.

b. $467 loss.

c. $1,000 gain.

d. $5,000 loss.
Answer #1

Value of $10000 bonds = 106000/100000*10000 = $10600

Premium = 10600 - 10000 = $600

Carrying value on May 1, 2018 = 10600 - 600*4/36 = $10533

Gain on retirement = 10000*1.1 - 10533 = $467

Option a. is correct answer.